# Grassroots Consensus

### v2

@@ 2.1 Blocklace Preliminaries
We use $a\neq b\in X$ as a shorthand for $a\in X\wedge b\in X\wedge a\neq b$. We assume a set of agents $\Pi$, each endowed with a unique key-pair and identified by its public key $p\in\Pi$. While $\Pi$ is potentially-infinite, we refer only to finite subsets $P\subset\Pi$ of it.
A $p$-block $b$, with $p\in\Pi$, is a triple $b=(h,x,H)$ where $h$ is the hash of the pair $(x,H)$ signed by $p$, referred to as the identifier of $b$, $x$ is an arbitrary payload, which may be empty, $x=\bot$, in which case we refer to $b$ as an empty block, and $H$ a finite set of block identifiers, the blocks of which are pointed to by $b$, with $b$ called initial if $H=\emptyset$. An encoding of such a triple $b=(h,x,H)$ via a sequence of bits in an agreed-upon form is referred to as a well-formed block. We assume the hash function is collision-free whp and cryptographic, so that blocks cannot form cycles and we may identify a block $b$ with its identifier $h$. Furthermore, the method of signature allows the public key $p$ to be recovered from a signature by $p$.
@@
The closure of a block $b$, denoted $[b]$, is the set of blocks observed by $b$. If a block variable $b$ is undefined then we deem its closure $[b]:=\emptyset$ to be empty rather than undefined. For a set of blocks $B$ the closure $[B]$ is defined by $[B]:=\bigcup_{b\in B}[b]$, which is also the set of blocks observed by $B$. A set of blocks $B$ is closed if $[b]\subseteq B$ for every $b\in B$, equivalently if $B=[B]$. A blocklace is a closed set of blocks.
To help avoid confusion, we use $B$ below to denote a closed set of blocks (a blocklace) and $D$ to denote any set of blocks, not necessarily closed. Given a set of blocks $D$, a pointer to a block $b$ is dangling in $D$ if $b\notin D$. Thus, a blocklace is a set of blocks with no dangling pointers. A block $b\in D$ is a tip of $D$ if no other block $b^{\prime}\neq b\in D$ observes $b$.
The depth $d(b)$ of a block $b$ is $1$ if $b$ is initial else $d(b)=d(b^{\prime})+1$ where $b^{\prime}$ is a maximal-depth block pointed to by $b$. The depth $d(B)$ of a set of blocks $B$ is defined by $d(B):=\max_{b\in B}d(b)$. A round $r\geq 1$ in $B$ is the set of all blocks $b\in B$ of depth $d(b)=r$, and the $r$-prefix of $B$, $B_{r}$, consists of all blocks $b\in B$ with depth $d(b)\leq r$, equivalently all rounds in $B$ up to and including $r$. Note that, by definition, all blocks of the same round do not observe each other.
@@ 2.2 Protocol Concepts
Models. We consider two models. The first is the standard model of eventual synchrony (aka partial synchrony) that assumes a reliable network in which every message sent among correct agents is eventually received. Message latency is unbounded, but in every run there is a time called the Global Stabilization Time (GST), which is unknown to the agents, after which all messages arrive within a least upper bound $\delta$ time for some $\delta>0$. In particular, all messages sent any time before GST arrive by time GST+$\delta$.
In the second model the assumption of a reliable network is relaxed. It assumes an eventually-reliable network, in which any message sent infinitely often between correct agents eventually arrives, and a message sent after GST between correct agents arrives within $\delta$. This is the model assumed in the original Paxos paper [26] and many sequels. We refer to this model as the UDP-ready model.
Parameters. The protocol employs three key parameters $(P,\sigma,\Delta)$:
@@
    $P\subset\Pi$, a set of agents,
@@
    $\frac{1}{2}\leq\sigma<1$, a fraction, used to specify a supermajority among the agents
  3. 3.
    $\Delta>0$, a presumed upper-bound on message delay after GST, known to the agents.
We refer to the parameters $(P,\sigma,\Delta)$ employed by the protocol for a given block $b$ or blocklace $B$, with values as above, as the prevailing constitution in $b$, resp. $B$; when the value of $\Delta$ is immaterial we may refer to the prevailing constitution $(P,\sigma)$. We begin by assuming the prevailing constitution to consist of arbitrary constant values for all blocks during the protocol run, as is usually the case. Later, in Section 6, we augment the protocol with constitutional amendment blocks by which the prevailing constitution may be amended.
Supermajorities. Given agents $P\subset\Pi$, $|P|=n$, and $\frac{1}{2}\leq\sigma<1$, a $\sigma$-supermajority among $P$ is a fraction $Q\subseteq P$ such that $|Q|>\sigma n$, or, equivalently, that $\frac{|Q|}{|P|}>\sigma$. A set of blocks $D$ is referred to as a $\sigma$-supermajority if there is a $\sigma$-supermajority $Q\subseteq P$ such that for every $q\in Q$, $D$ includes at least one $q$-block.
Typically, a permissioned consensus protocol assumes that there are at most $f<n$ faulty agents among $P$. Given $f$, we say that a blocklace $B$ is $f$-safe if it includes equivocations by at most $f$ agents, a block $b$ is $f$-safe if $[b]$ is $f$-safe, and define $\sigma:=\frac{n+f}{2n}$. For example, if $f=0$ then $\sigma=\frac{1}{2}$; if $f=\frac{1}{2}n$ then $\sigma=\frac{3}{4}$; if $n=3f+1$ (the standard assumption), then $\sigma=\frac{n+\frac{n-1}{3}}{2n}=\frac{2}{3}-\frac{1}{6n}$, namely a supermajority greater-or-equal to $\frac{2}{3}n$ is required.
Under this assumption, a $\sigma$-supermajority includes a majority of the correct agents, and hence the intersection of any two $\sigma$-supermajorities includes at least one correct agent. Typically, permissioned consensus protocols refer to “blocks by at least $n-f$ agents among $P$”, which is equivalent to “a $\sigma$-supermajority among $P$” in the standard case of $n=3f+1$.
Waves. A blocklace is seen as a sequence of waves, where the $k^{th}$-wave in $B$ , $k\geq 1$, consists of three consecutive rounds $r$, $r+1$, $r+2$ in $B$, $r=3(k-1)+1$, referred to as the first, second, and third round of the wave.
For a given $P\subset\Pi$, the partial function leaderP assigns a leader $p\in P$ for the first round of every wave, where a first-round $p$-block $b$ with prevailing constitution $(P,\sigma)$ is a leader block if $p=\textit{leader}_{P}(d(b))$. We assume the leader function to be fair in that for every wave and any $p\in P$ there is a subsequent wave for which $p$ is the leader of its first round.
The consensus protocol progresses in waves. In each wave, a first round block is a candidate for finality in two cases—if it is a leader block or the only block in its round. These two possibilities typically arise in high- and low-throughput scenarios, respectively. During low throughput, agents quiesce, not producing a first-round block until they have a new transaction; and when an agent produces a first-round block $b$ that does not conflict with others, all agents produce second and third round blocks that finalize $b$. During high throughput, conflicting blocks may be produced and the leader function is used to break symmetry; leader blocks are finalized and induce the total ordering of all blocks.
A second-round block $b$ of a wave may approve any number of first-round blocks, but endorses (towards finality) at most one of them, defined as follows. Note that the next two definitions are mutually-recursive:
Definition 2.1 (Endorse).
Let $b$ be a first-round block with prevailing constitution $(P,\sigma)$ and $b^{\prime}$ a second-round block of the same wave in $B$. Then $b^{\prime}$ endorses $b$ if the previous wave is:
@@
    quiescent and $b$ is the only block in its round approved by $b^{\prime}$, or
@@
    non-quiescent and $b$ is a leader block approved by $b^{\prime}$.
A set of blocks $D$ endorses $b$ if every $b^{\prime}\in D$ endorses $b$.
A first-round block $b$ is final if the third round of the same wave has a $\sigma$-supermajority that approves a second-round $\sigma$-supermajority that endorses $b$. Formally (text in parenthesis is explanatory, the definition is valid also without it):
Definition 2.2 (Ratified, Final, Ordered, Quiescent).
Given a blocklace $B$, a block $b\in B$ with prevailing constitution $(P,\sigma)$ is:
@@
    ratified by a block $b^{\prime}$ if $b^{\prime}$ approves a $\sigma$-supermajority that endorses $b$;
@@
    final in $B$ if there is a $\sigma$-supermajority in the third-round of $b$’s wave in $B$, in which each block ratifies $b$, referred to as a finalizing $\sigma$-supermajority.
A wave in such a blocklace $B$ is:
  1. 1.
    finalizing if it includes a (first-round) final block $b$; and
  2. 2.
    quiescent if in addition all the wave’s blocks except perhaps $b$ are empty and $b$ does not conflict with any block in $B$.
As an initial condition, we refer to the (non-existent) wave that ends in round 00 as quiescent. A block $b$ is slow if it is both finalized by a quiescent wave and follows a quiescent wave.
Note that since only first-round blocks are endorsed, they are the only ones that can be ratified or final. We observe that the block properties defined above are monotonic: {observation}[Monotonicity] If a block is ratified or final in some blocklace $B$, it is also ratified or final, respectively, in any blocklace $B^{\prime}\supset B$. Note, however, that quiescence is not monotonic, both because conflicting blocks may be added to a round, and because more rounds can be added to the blocklace.
@@
Quiescence, low- and high-throughput scenarios. The protocol is initially quiescent. It begins operations and continues to do so as long as agents have payloads to send. If multiple agents send payloads in the same wave, the protocol switches to high-throughput mode. It returns to being quiescent following a quiescent wave.
During low-throughput, in the good-case scenario, all waves are quiescent. A wave commences when some agent has a new transaction. This agent issues a (first-round) non-empty block, and all agents follow by issuing two rounds of blocks: Second-round blocks that endorse the first-round block, and third-round blocks each approving all second-round blocks, upon which the first-round block becomes final. The new wave is also quiescent provided all second- and third-round blocks are empty. The agent that issued a finalized, non-empty, first-round block in this scenario is referred to as a spontaneous leader, in contrast with the formal leader assigned by the leaderP function.
During high-throughput, no wave is quiescent, and the good-case scenario unrolls as above, except that the first-round block is issued promptly by the wave’s (formal) leader, whether or not it has any payload, and second- and third-round blocks are typically non-empty (contain transactions to be ordered by the next-wave’s final leader block). The good-case scenarios are summarized in Figure 1.
Going beyond the good case, if during high-throughput agents fail to receive a first-round leader block in a timely manner, they issue first-round blocks of their own. Once a super-majority of agents do so, agents can continue to the next round without waiting for a leader block. And if during low-throughput there are multiple agents issuing first-round blocks, then it is possible that none of the first round blocks will become final. In this case, the wave is not quiescent and the protocol proceeds as in the high-throughput scenario.
@@
Given a blocklace $B$ with prevailing constitution $(P,\sigma)$, a round $r>0$ in $B$ is advanced if $r$ is a:
@@
    first round that either (i) includes a $\sigma$-supermajority of blocks among $P$, or (ii) includes a leader block or (iii) follows a quiescent wave and includes at least one block, or a
@@
    second or third round that includes a $\sigma$-supermajority of blocks among $P$.
Round 00 is trivially advanced in all blocklaces.
@@
Agents incorporate blocks in the blocklace only if they are valid, namely could have been produced by a correct agent during the run of the protocol. Formally, an ordinary block $b$ is valid if it is well-formed and round $d(b)-1$ is advanced in $[b]$.
Round $r+1$ has a single leader block (red), round $r+2$ consists of a supermajority of blocks (blue as an example) pointing only to (and thus endorsing) the first-round leader block, and round $r+3$ consists of a supermajority of blocks (green as an example) each pointing to (and thus approving) all and only the second-round blocks, and thus ratifying the first-round leader block.
In the low-throughput scenario, the previous wave ending in round $r$ is quiescent, the spontaneous leader block (red) may be produced at any time after the previous round $r$ has advanced, and all other blocks are empty, thus the wave $w$ is also quiescent.
In the high-throughput scenario, the formal leader produces a block (red) within $\delta$ after the previous round $r$ is advanced, and all (or many) of the blocks are non-empty.
Figure 1: A good-case wave.
Dissemination. The protocol uses three types of dissemination-inducing blocks: ack, nack, and nudge, defined below; other blocks are referred to as ordinary.
@@
Dissemination-inducing blocks are not incorporated in the blocklace of their sender or recipient. Hence, they contribute only to the liveness of the protocol and do not affect its safety, which is derived from structural properties of the blocklace.
@@
The Grassroots Consensus protocol we present in Section 2.3 assumes reliable links and uses nack and nudge blocks for dissemination. Unlike Cordial Miners, which incorporates a blocklace-based dissemination protocol with $O(n^{2})$ amortized communication complexity in the good case, Grassroots Consensus for this model achieves $O(n)$ amortized communication complexity in the good case during high-throughput. The UDP-ready protocol (given in Section 4) uses also ack blocks for dissemination.
@@
nack blocks are used to achieve reliable broadcast of ordinary blocks. They are needed in order to address partial dissemination by Byzantine agents. For example, consider a Byzantine agent that sends a block $b$ to correct agent $p$ but not to another correct agent $q$, and $p$ adds $b$ into its blocklace and sends a block $b^{\prime}$ that depends on $b$. Then when $q$ receives $b^{\prime}$, it cannot add it to its blocklace because its predecessor $b$ is missing. To this end, $q$ sends a nack block to $p$ specifying the blocks it is missing.
@@
Note that out-of-order blocks may naturally arise on a reliable network even during periods of synchrony, even when there are no Byzantine agents, because there is no bound on how quickly messages can arrive — in the example above, $b$’s sender might be correct and $b$ may still be under way to $q$. We therefore introduce a delay before sending a nack. Thus, in the good case, when the network is synchronous and there is no partial-dissemination by Byzantine agents, every correct agent sends its newly created blocks to all other agents and no further communication is needed.
@@
The following definition relates to a setting in which an agent $p$ has a blocklace $B$ and it receives a new block $b$. If $[b]\not\subseteq B\cup\{b\}$, then $p$ cannot incorporate $b$ into $B$ since some pointers in $b$ are dangling in $B$. While obviously $p$ cannot know all the blocks it does not know, it can know the boundary of its knowledge—where $p$’s ignorance starts in relation to $b$. These are the tips of $[b]\setminus(B\cup\{b\})$. Note that if no block pointed to by $b$ is in $B$, then the pointers in $b$ point exactly to tips in $[b]\setminus(B\cup\{b\})$. Hence the following definition:
@@
Given a blocklace $B$, a nack-block for a block $b=(h,x,H)$ is the block $b^{\prime}=(h^{\prime},x^{\prime},H^{\prime})$ where $x^{\prime}=(\textsc{nack},h)$ and $H^{\prime}\neq\emptyset$ contains the tips of $[b]\setminus(B\cup\{b\})$.
Note that as defined above, $[b^{\prime}]\subseteq[b]$. Also note that, given a block $b=(h,x,H)$ and a blocklace $B$, the tips of $[b]\setminus(B\cup\{b\})$ are observed directly from $b$: These are the pointers in $H$ that are dangling in $B$, namely point to blocks not in $B$. These dangling pointers attest to the ‘knowledge gap’ of the recipient of $b$, namely, which blocks need to be added to $B$ so that $b$ will be in order following $B$. If there is no such gap, namely $[b]\subseteq B\cup\{b\}$, then the nack-block for $b$ and $B$ is undefined.
In the reliable network model, the reason for the asymmetry of having nack-blocks but not ack-blocks is that any ordinary $p$-block is in fact an ack-block, as it attests to all the blocks known to $p$ at the time of its creation; ack-blocks will be added in the UDP-ready protocol described in Section 4, when eventually-reliable networks are considered.
@@
nack blocks enable pull-based forwarding, where an agent who knows it misses a block asks for it. In addition, we make use of push-based forwarding to facilitate progress using nudge-blocks.
Definition 2.5 (nudge-Block).
@@
Given a blocklace $B$, a nudge-block $(h,x,H)$ for round $r+1>1$ has payload $x=\textsc{nudge}$ and pointers $H$ to the blocks of depth $r$ in $B$.
@@
If an agent $p$ suspects that the leader of the next round $r+1$ does not proceed since it is missing some round $r$ blocks, $p$ sends the leader a nudge-block pointing to the $r$ blocks in its blocklace. If the leader indeed misses any of these blocks, which after GST may happen due to Byzantine partial dissemination, it will respond to $p$ with a nack-block, which will in turn cause $p$ to send to the leader the blocks it needs.
2.3 The Protocol
Here we present Grassroots Consensus for eventual synchrony with a reliable network. We prove the protocol’s safety and liveness in Section 3, and present its UDP-ready extension in Section 4. In this section, we present the protocol using arbitrarily-fixed values for the parameters $(P,\sigma,\Delta)$. These can be changed using constitutional amendments, discussed in Section 6.

### v4

@@ 3.1 Blocklace Preliminaries
We use $a\neq b\in X$ as a shorthand for $a\in X\wedge b\in X\wedge a\neq b$.
A $p$-block $b$, with $p\in\Pi$, is a triple $b=(h,x,H)$ where $h$ is the hash of the pair $(x,H)$ signed by $p$, referred to as the identifier of $b$, $x$ is an arbitrary payload, which may be empty, $x=\bot$, in which case we refer to $b$ as an empty block, and $H$ a finite set of block identifiers, the blocks of which are pointed to by $b$. A block with no pointers, $H=\emptyset$, is called initial, in which case it can also be written as a pair $b=(h,x)$. An encoding of a block via a sequence of bits in an agreed-upon form is referred to as a well-formed block. We assume the hash function to be collision-free whp and cryptographic, so that blocks cannot form cycles and we may identify a block $b$ with its identifier $h$. Furthermore, the method of signature allows the public key $p$ to be recovered from a signature by $p$.
A set of blocks $X$ is referred to as a $\sigma$-supermajority given a set of agents $P$ if there is a $\sigma$-supermajority $Q\subseteq P$ such that for every $q\in Q$, $X$ includes at least one $q$-block.
@@
The closure of a block $b$, denoted $[b]$, is the set of blocks observed by $b$. For a set of blocks $B$, its closure $[B]$ is defined by $B]:=\bigcup_{b\in B}[b]$, which is also the set of blocks observed by $B$. A set of blocks $B$ is closed if $b]\subseteq B$ for every $b\in B$, equivalently if $B=[B]$. A blocklace is a closed set of blocks.
To help avoid confusion, we use $B$ to denote a closed set of blocks (a blocklace) and $D$ or $X$ to denote any set of blocks, not necessarily closed. Also, we refer to a set of blocks by agents in $P\subset\Pi$ as $P$-blocks, the let the $P$-blocks in $B$ denote the set of all blocks in $B$ by agents in $P$.
Given a set of blocks $D$, a pointer to a block $b$ is dangling in $D$ if $b\notin D$. Thus, a blocklace is a set of blocks with no dangling pointers. A block $b\in D$ is a tip of $D$ if no other block $b^{\prime}\neq b\in D$ observes $b$.
The depth of a block $b$, $\textit{depth}(b)$, is 0 if $b$ is initial else $\textit{depth}(b)=\textit{depth}(b^{\prime})+1$ where $b^{\prime}$ is a maximal-depth block pointed to by $b$. The depth of a set of blocks $B$ is defined by $\textit{depth}(B):=\max_{b\in B}\textit{depth}(b)$. In the context of blocks, we use most-recent as synonym for deepest. A round $r\geq 0$ in $B$ is the set of all blocks $b\in B$ with $\textit{depth}(b)=r$, and the $r$-prefix of $B$, $B_{r}$, consists of all blocks $b\in B$ with $\textit{depth}(b)\leq r$, equivalently all rounds in $B$ up to and including $r$. Note that, by definition, blocks of the same round do not observe each other.
@@ 3.2 Protocol Concepts
The blocklace created during an epoch begins with a constitutional genesis block $(h,d)$ where $d$ specifies a new constitution $C=(P,\sigma,\Delta)$. In fact $d$ also records the constitutional decision to amend the old constitution (if any) to the new constitution and is signed by multiple agents; these are discussed in Section 5. Here, we focus on the behaviour of the protocol during a single epoch. For now we refer to a blocklace in which all initial blocks are constitutional genesis blocks as a constitutional blocklace. Every participant $p\in P$ commences an epoch $b$ with its blocklace including $b$ and its local epoch identifier $e$ set to $b$, $e\leftarrow b$.
The protocol employs an evolving shared blocklace. Each participant $p\in P$ maintains a local copy of the constitutional blocklace. In the course of an epoch, agents create new blocks that observe the epoch’s genesis block and and disseminate them to other agents. Blocks may include transactions the agents wish to order, if they have any. While a blocklace is a partially ordered set of blocks, the protocol totally orders the blocks in an epoch in a consistent way, producing the same output sequence at all participants in an epoch. Moreover, epochs are also totally ordered, so that the totality of all blocks of all epochs are also totally ordered, even though not always by the same participants
Because correct agents follow the protocol, they do not equivocate. Hence, for a blocklace $B$ created by the protocol with constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents.
Waves. Given a constitutional blocklace $B$, a wave $W\subseteq B$ in a blocklace is either its constitutional genesis block, referred to as its $0^{th}$-wave, or the set of blocks of depths $r+1$, $r+2$, and $r+3$ in $B$, referred to as the the first, second, and third round of the wave, respectively, for any $r\geq 0$ such that $r\mod 3=0$. Such a wave is the $k^{th}$ wave of $B$, $k:=(r+3)/3$.
For a given $P\subset\Pi$, the function leaderP maps integers to $P$. For each first round of a wave it assigns a formal leader $p\in P$, where a first-round $p$-block $b$ is a leader block if $p=\textit{leader}_{P}(\textit{depth}(b))$. We assume the leader function is defined so that it is fair, in that for every wave and every $p\in P$ there is a subsequent wave for which $p$ is the formal leader of its first round.
Finalizing blocks in waves. The consensus protocol commences with a constitutional blocklace and progresses in waves. The protocol strives to finalize a new block in each wave. The first round of a wave includes candidates for finality, the second round includes endorsements of candidates, and the third round includes ratifications.
A first round block is a candidate for finality in two cases—if it is a formal leader block or the only block in its round. These two possibilities typically arise in high- and low-throughput scenarios, respectively. During low throughput, agents quiesce, not producing a first-round block until they have a new transaction; and when an agent produces a first-round block $b$ that does not conflict with any others, all agents produce second and third round blocks that finalize $b$, in which case $b$ is referred to as spontaneous leader. During high throughput, multiple agents may have transactions to send, and the leader function is used to break symmetry, so that only the formal leader sends a leader block in the first round. Finalized leader blocks of both types induce the total ordering of all blocks, as in other DAG-based protocols [17, 18, 23].
A second-round block $b$ of a wave may approve any number of first-round blocks, but endorses (towards finality) at most one of them, defined as follows. Note that the next two definitions are mutually-recursive.
Definition 3.1 (Endorse).
Let $b^{\prime}$ be a second round block that approves a first round block $b$ in a wave $W$ in the blocklace $B=[b^{\prime}]$. Let $(P,\sigma)$ be the prevailing constitution.
Then $b^{\prime}$ endorses $b$ if the most-recent wave $W^{\prime}\subset B$ that precedes $b$ is:
@@
    quiescent and $b$ is the only first-round block approved by $b^{\prime}$, or
@@
    non-quiescent and $b$ is a $\textit{leader}_{P}$ block approved by $b^{\prime}$.
A set of second-round blocks $X\subset W$ endorses $b$ if every $b^{\prime}\in X$ endorses $b$.
Note that endorsement is an invariant block property, as it depends only on $b^{\prime}]$.
Next, we specify the finality rule. A first-round block $b\in W$ is final if the third round of $W$ has a $\sigma$-supermajority that approves a second-round $\sigma$-supermajority that endorses $b$. Formally:
Definition 3.2 (Ratified, Final, Ordered, Quiescent).
Assume a blocklace $B$ and a wave $W\subseteq B$. A first-round block $b\in W$ with prevailing constitution $(P,\sigma)$ is:
@@
    ratified by a third-round block $b^{\prime}\in W$ if $b^{\prime}$ approves a $\sigma$-supermajority in $P$ that endorses $b$, in which case $b$ is also ratified in $W$ and in $B$; and
@@
    final in $W$ and in $B$ if the third-round of $W$ includes a $\sigma$-supermajority in $P$, referred to as a finalising $\sigma$-supermajority, in which each block ratifies $b$.
The wave $W$ is:
@@
    finalising if it includes a block $b$ that is final in $W$; and
@@
    quiescent if, in addition, all blocks in $W$ except perhaps $b$ are empty, and no block in $B$ conflicts with $b$.
In addition, the $0^{th}$-wave is defined to be quiescent and its constitutional genesis block to be final.
Note that since only first-round blocks are endorsed, they are the only ones that can be ratified or final. Notice our use of the term “finalised” only applies to such leader blocks; the finalisation of leader blocks induces an order on all the blocks they approve, which are then said to be ordered.
We observe that the block properties defined above are monotonic: {observation}[Monotonicity] If a block is ratified or final in some blocklace $B$, it is also ratified or final, respectively, in any blocklace $B^{\prime}\supset B$. Note, however, that quiescence is not monotonic, because conflicting and non-empty blocks may be added to a round.
@@
Proof 3.3.
The observation follows from the fact that a second-round block endorses at most one first-round block in any blocklace it is included in, so to have two $\sigma$-supermajorities of the same round endorse two blocks, at least one correct agent must endorse two blocks in this round; to do so this correct agent must equivocate, a contradiction.
Quiescence, low- and high-throughput scenarios. The protocol is initially quiescent. It begins operations and continues to operate as long as agents have payloads to send. If multiple agents send payloads in the same wave, the protocol switches to high-throughput mode. It returns to being quiescent following a quiescent wave.
Round $r+1$ has a single leader block (red), round $r+2$ consists of a supermajority of blocks (blue as an example) pointing only to (and thus endorsing) the first-round leader block, and round $r+3$ consists of a supermajority of blocks (green as an example) each pointing to (and thus approving) all and only the second-round blocks, and thus ratifying the first-round leader block, which is thus finalized by the wave.
Low-throughput good case: The previous wave ending in round $r$ is quiescent, the spontaneous non-empty leader block (red) may be produced at any time after the previous round $r$ has advanced, followed promptly by two rounds of empty blocks, resulting in a finalising and quiescent wave.
High-throughput good case: The formal leader produces a block (red) within $\delta$ after the previous round $r$ is advanced and all (or many) of the wave’s blocks are non-empty.
Figure 1: A good-case wave $w$.
We first consider the good-case scenarios summarized in Figure 1. During low-throughput all waves are quiescent. A wave commences when some agent has a new transaction. This agent issues a (first-round) non-empty block, and all agents follow promptly with two rounds of empty blocks: Second-round blocks that endorse the first-round block, and third-round blocks each approving all second-round blocks, upon which the first-round block becomes final, with the new wave also being quiescent.
During high-throughput, no wave is quiescent, and the good-case scenario unrolls as above except that the first-round block is issued promptly by the wave’s (formal) leader, whether or not it has any payload, and second- and third-round blocks are typically non-empty, with transactions to be ordered by the the next final leader block.
Going beyond the good case, if during high-throughput agents fail to receive a first-round leader block in a timely manner, they issue first-round blocks of their own. Once a super-majority of agents do so, agents can continue to the next round without waiting for a leader block, with blocks by correct agents to be ordered by the leader of some subsequent finalising wave. Similarly, if during low-throughput there are multiple agents issuing first-round blocks, then typically none of the first-round blocks will become final. In this case, the wave is not quiescent and the protocol proceeds as in the high-throughput scenario.
@@
Assume a constitutional blocklace $B$ with prevailing constitution $(P,\sigma)$. A round $r$ in a wave $W\subseteq B$ is advanced if it is a:
@@
    first round that either (i) includes a $\sigma$-supermajority of blocks among $P$, or (ii) includes a formal leader block or (iii) is preceded by a quiescent wave in $B$ and includes at least one block, or a
@@
    second or third round that includes blocks by a $\sigma$-supermajority among $P$.
Round 0 is trivially advanced.
@@
Agents incorporate blocks in the blocklace only if they are valid, namely could have been produced by a correct agent during the run of the protocol. Formally:
Definition 3.5 (Valid Block).
An ordinary block $b$ is valid if it is well-formed and round $\textit{depth}(b)-1$ is advanced in $[b]$.
@@
Dissemination. The protocol uses two types of dissemination-inducing blocks: nack and inform; other blocks are referred to as ordinary. Dissemination-inducing blocks are not incorporated in the blocklace of their sender or recipient. Hence, they contribute only to the liveness of the protocol and do not affect its safety, which is derived from structural properties of the blocklace.
@@
Unlike Cordial Miners, which incorporates a blocklace-based dissemination protocol with $O(n^{2})$ amortized communication complexity in the good case, Constitutional Consensus achieves for this model $O(n)$ amortized communication complexity in the good case during high-throughput.
Protocols such as TCP achieve the reliable transfer of a sequence of bytes from the sender to the recipient by employing ack and nack messages for transmission control. The purpose of an ack message is to acknowledge the receipt of a datagram, and of a nack message to report an out-of-order datagram to its sender, e.g., in the case of receiving the $(k+1)^{th}$-datagram before receiving the $k^{th}$-datagram.
In the context of the blocklace, an ‘out-of-order’ block is a block received before some block it points to, and the purpose of a nack-block is indeed to report of such an out-of-order block to its creator (which is not necessarily its sender).
@@
nack blocks are used to achieve reliable broadcast of ordinary blocks. They are needed in order to rectify partial dissemination by Byzantine agents. For example, consider a Byzantine agent that sends a block $b$ to correct agent $p$ but not to another correct agent $q$, and $p$ adds $b$ into its blocklace and sends a block $b^{\prime}$ that depends on $b$. Then when $q$ receives $b^{\prime}$, it cannot add it to its blocklace because its predecessor $b$ is missing. To this end, $q$ sends a nack block to $p$ specifying the blocks it is missing.
@@
Note that out-of-order blocks may naturally arise on a reliable network even during periods of synchrony and even when there are no Byzantine agents, because there is no bound on how quickly messages can arrive — in the example above, $b$’s sender might be correct and $b$ may still be under way to $q$. We therefore introduce a delay before sending a nack. The result is that in the good case, when the network is synchronous and there is no partial-dissemination by Byzantine agents, every correct agent sends its newly created blocks to all other agents and no nack-blocks or other communication is needed.
@@
The following definition relates to a setting in which an agent $p$ has a blocklace $B$ and it receives a new block $b$. If $b]\not\subseteq B\cup\{b\}$, then $p$ cannot incorporate $b$ into $B$ since some pointers in $b$ are dangling in $B$. These are the tips of $b]\setminus(B\cup\{b\})$. Hence the following definition:
@@
Given a blocklace $B$, a nack-block by $p$ for a block $b=(h,x,H)$ such that $b]\not\subseteq B\cup\{b\}$ is the $p$-block $b^{\prime}=(h^{\prime},x^{\prime},H^{\prime})$ where $x^{\prime}=(\textsc{nack},h)$ and $H^{\prime}\neq\emptyset$ contains the tips of $b]\setminus(B\cup\{b\})$.
Note that as defined above, $b^{\prime}]\subseteq[b]$, and if $b$ has no pointers dangling in $B$ then a nack-block for $b$ and $B$ is undefined. If the nack-block $b^{\prime}$ is defined, then its pointers $H^{\prime}$ are precisely the non-empty subset of $H$ that is dangling in $B$.
@@
The reason for the asymmetry of having nack-blocks but not ack-blocks is that any ordinary $p$-block functions as an ack-block, as it attests to all the blocks known to $p$ at the time of its creation. nack blocks enable pull-based forwarding, where an agent who knows it misses a block asks for it. In addition, we make use of push-based forwarding to facilitate leader progress using inform-blocks.
Definition 3.7 (inform-Block).
@@
Given a blocklace $B$, an inform-block $(h,x,H)$ for round $r+1>1$ has payload $x=\textsc{inform}$ and pointers $H$ to the blocks of depth $r$ in $B$.
@@
If an agent $p$ suspects that the leader of the next round $r+1$ does not proceed since it is missing some round $r$ blocks known to $p$, then $p$ sends the leader an inform-block pointing to the $r$ blocks in its blocklace. If the leader indeed misses any of these blocks, which may happen even after GST due to Byzantine partial dissemination, it will respond to $p$ with a nack-block, which will in turn cause $p$ to send to the leader the blocks it needs.
3.3 Single-Epoch Constitutional Consensus Pseudocode